# DiffusionAvatars: Deferred Diffusion for High-fidelity 3D Head Avatars

### v1

@@ Abstract
We train DiffusionAvatars on RGB videos and corresponding tracked NPHM meshes of a person and test the obtained avatars in both self-reenactment and animation scenarios. Our experiments demonstrate that DiffusionAvatars generates temporally consistent and visually appealing videos for novel poses and expressions of a person, outperforming existing approaches.
(a) Avatar creation
(b) Avatar animation
Figure 1: Given a set of multi-view videos and corresponding tracked meshes, we build a DiffusionAvatar of a person. Our method translates expressions of a morphable model into realistic facial appearances of a person while also providing control over the viewpoint.
@@ 1 Introduction
A significant problem in computer vision and graphics is the creation of photorealistic animatable avatars. Ideally, these avatars permit consistent video renderings while providing free control over desired expression, pose, and viewpoint. A digital 3D head avatar may then be used in various scenarios such as VR/AR applications, immersive teleconferencing, gaming, movie animation, language learning, or as a virtual assistant.
@@ 3.2 NPHM
NPHM [17] is a morphable head model that generates a signed distance field (SDF) of a human head given an identity code $z_{id}$ and an expression code $z_{exp}$. We employ COLMAP [59, 58] to obtain pointclouds $\{\mathcal{P}_{t}\}$ for each timestep $t$ of a multi-view video sequence of the person. Subsequently, we fit NPHM to each pointcloud $\mathcal{P}_{t}$ to obtain $z_{id}$ and $z^{t}_{exp}$. We use a variant of NPHM that uses a backward deformation field instead of the originally proposed forward deformation field as it simplifies the fitting process. Formally, we obtain the NPHM fitting as follows:
@@ 3.6 Deferred Diffusion
NeRFace [15]
DNR [68]
DNR + GAN
DiffusionRig [11]
Ours
GT
Figure 3: Qualitative results for self-reenactment. We compare against 3D methods (NeRFace [15]) and methods employing 2D renderers (Deferred Neural Rendering [68], DiffusionRig [11]). Note that the slightly gray background for DiffusionRig is caused by their training scheme using $\epsilon$-prediction (see Sec. 3.1). Our method consistently produces more expressive facial performances while simultaneously providing more detailed renderings.
Source Actor
NeRFace [15]
DNR [68]
DNR + GAN
DiffusionRig [11]
Ours
@@ 4.1 Training
For our experiments, we employ the recently released multi-view video dataset from NeRSemble [33]. The dataset contains 26 different multi-view sequences per person captured with 16 cameras, featuring different facial expressions, sentences and emotions. We use BackgroundMattingV2 [38] to remove the background and a segmentation network to remove the torso [86]. We fit NPHM on the video sequences of 5 different persons to create pairings of multi-view videos and tracked meshes. In total, we use about 3300 timesteps per person. We train one DiffusionAvatar for each person on these pairings with Adam [32] using a learning rate of $41e-$ 4 for ControlNet and the expression conditioning layers, and a learning rate of $21e-$ 2 for the TriPlane feature lookups. We use a batch size of $B=8$ and spatial resolutions $H=W=512$. We employ random cropping and resizing of the target images to avoid overfitting. The size of our spatial feature maps is $512\times 512\times 16$ for each TriPlane and the ambient map. In total, this yields 73 channels for the input renderings $R_{t}$ fed to our diffusion-based renderer. The pre-trained LDM is a Stable Diffusion v2.1 network [55]. Training takes roughly two days on a single RTX A6000 for 100k gradient descent steps.
@@ 4.2 Experiment Setup
We measure the quality of a 3D avatar in terms of image quality, 3D consistency and expression synthesis. To this end, we conduct experiments on two tasks: Self-reenactment and avatar animation. For self-reenactment, we hold out 2 complete sequences and the frontal camera of every sequence during training to measure both image quality and view consistency. During testing, we provide the holdout tracked meshes of the same person and render from the novel view. For avatar animation, we first generate NPHM meshes using the same identity code $z_{id}$ as the avatar and a sequence of expression codes $\{z^{t}_{exp}\}$ from a different person. We then render a video of the avatar from a novel view. To evaluate the realism of this expression transfer, we conduct a user study where we show the animated avatars to participants.
@@
We employ two paired-image metrics to measure the quality of individual generated frames: Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [85]. Furthermore, to assess temporal consistency, we employ the perceptual video metric JOD [43] to measure visual similarity of a rendered video to its ground truth counterpart. Finally, we make use of two face-specific metrics: Average Keypoint Distance (AKD) measured in pixels with keypoints estimated from PIPNet [28], and cosine similarity (CSIM) of identity embeddings based on ArcFace [9].
@@
Deferred Neural Rendering (DNR) [68]. A screen-space decoder built on top of Pix2Pix [89] that uses learnable neural textures rigged to a tracked Basel Face Model [51]. We provide our tracked NPHM meshes as input for better comparison. Additionally, we compare against a version of DNR that also uses an adversarial loss for sharper images that we call DNR+GAN in the following.
@@ 4.3 Self-Reenactment
In Tab. 1, we compare DiffusionAvatars to other state-of-the-art approaches on the self-reenactment task. Results are averaged over avatars generated from five different persons. Our method shows the overall best performance with the biggest improvement in perceptual and face-specific metrics. The results indicate that DiffusionAvatars generates sharper images (LPIPS) that are more temporally consistent (JOD). Furthermore, our method better maintains the person’s identity when rendered from a novel view (CSIM) than the baselines. Finally, facial features such as the position of mouth corners and eyebrows synthesized by our method match the ground truth expressions more closely (AKD).
@@
Method Self Reenactment User Study
@@
NeRFace [15] 22.5 0.299 6.65 5.10 0.761 2.19 2.35
DiffusionRig [11] 19.6 0.218 6.42 2.66 0.886 2.47 3.52
DNR [68] 24.4 0.226 7.31 2.18 0.898 2.87 3.94
DNR+GAN [68] 22.8 0.115 7.06 2.32 0.859 3.06 3.97
Ours 24.5 0.088 7.53 1.86 0.912 4.02 4.14
Table 1: Quantitative Comparison: We report metrics for a self-reenactment scenario on unseen expressions and unseen views. For our user study on avatar animation, we collect 735 responses from 49 participants to measure Visual Quality (VQ) and Driving Fidelity (DF) on a scale of 1-5.
@@ 4.4 Avatar Animation
Due to the task’s inherent lack of any ground truth, we perform a user study to validate avatar animation. We collect 735 responses from 49 participants on two questions: (i) Visual Quality (VQ): What is the overall quality of the avatar animation? (ii) Driving Fidelity (DF): How closely does the animated avatar match the expressions of the driving sequence? We present two videos side-by-side: One sequence of a source actor performing various expressions, and a video of another person’s avatar with the expressions transferred to it. The results of our user study can be seen in Tab. 1. Our method achieves the highest scores for both visual quality and similarity to the driving sequence. The difference is particularly pronounced for the visual quality score. We attribute this to the fact that users heavily penalize any visual artifacts immediately obvious in video renderings.
These perceptual differences in the task of avatar animation are also apparent in the qualitative comparison shown in Fig. 4.
@@ 4.5 Ablations
In the following, we examine the design choices of our method. All ablation experiments are conducted on 3 avatars on the self-reenactment task, and the averaged results are summarized in Tab. 2.
@@
Table 2: Ablation of architectural choices. Our full method utilizing TriPlane feature mapping performs overall the best. The biggest quantitative improvements stem from using NPHM instead of FLAME, employing direct expression conditioning, and using a pre-trained LDM.
@@
We study whether diffusion models are necessary for our task. To this end, we train our architecture without the pre-trained LDM to directly predict the output in a single forward pass, similar to Deferred Neural Rendering. The results indicate that such a method performs notably worse, especially in the LPIPS metric, indicating that diffusion is important to obtain sharp renderings.
@@
We compare DiffusionAvatars with a version trained on renderings of a fitted FLAME instead of NPHM. Since NPHM models facial geometry as an SDF, it possesses a much higher representational power than FLAME. As a result, the renderings are much closer to the actual desired appearance of the head. This is also reflected by the results of our ablation study, where using FLAME gives consistently worse performance. This is especially true for the AKD metric, indicating that our diffusion-based neural renderer cannot reenact certain complex expressions based on FLAME alone.
@@
The overall performance worsens when direct expression conditioning is turned off. We attribute this to the fact that expression conditioning helps to distinguish subtle expression details that can hardly be inferred from the NPHM mesh alone. It also helps DiffusionAvatars to synthesize expressions not captured by NPHM, such as tongue movement, by enabling the network to correlate such appearances with certain expression code combinations.
@@
We compare three spatial feature mapping approaches: (i) Our Triplanes, (ii) a simpler approach based on spherical feature mapping, and (iii) not using any spatial features at all. For spherical feature mapping, we create an approximate UV map for each NPHM mesh by projecting the 3D canonical coordinates for each vertex onto a sphere centered inside the head. Tab. 2 shows that using TriPlanes boosts most metrics, whereas the simpler spherical mapping brings almost no benefit.
@@ Acknowledgements
Appendix A Novel View Synthesis Results
@@
Table 3: Quantitative Comparison: We report metrics for novel-view-synthesis.
@@ Appendix B User Study Setup
Appendix C Qualitative Comparison of Ablations
Appendix D Societal Impact
DiffusionAvatars provides the means to generate realistic images of faces with control over viewpoint, pose, and expression. As such, there is a risk of misusing an avatar to the original person’s disadvantage. However, we focus on 3D applications such as immersive teleconferencing or character animation, not indistinguishable 2D video synthesis. Our method does not model lighting, the background, or the torso. Therefore, we believe a detection system can be built to identify generations of our method by exploiting such characteristics [56].
@@ Appendix D Societal Impact
FLAME [35]
Ours
GT
w/o exp. cond.
Ours
GT
w/o LDM prior
Ours
GT

### v2

@@ Abstract
We train DiffusionAvatars on RGB videos and corresponding fitted NPHM meshes of a person and test the obtained avatars in both self-reenactment and animation scenarios. Our experiments demonstrate that DiffusionAvatars generates temporally consistent and visually appealing videos for novel poses and expressions of a person, outperforming existing approaches.
(a) Avatar creation (b) Avatar animation
Figure 1: Given a set of multi-view videos and corresponding fitted meshes, we build a DiffusionAvatar of a person. Our method translates expressions of a morphable model into realistic facial appearances of a person while also providing control over the viewpoint. Project website: https://tobias-kirschstein.github.io/diffusion-avatars/
@@ 1 Introduction
A significant problem in computer vision and graphics is the creation of photorealistic animatable avatars. Ideally, these avatars permit consistent video renderings while providing free control over desired expression, pose, and viewpoint. A digital 3D head avatar may then be used in various scenarios such as VR/AR applications, immersive teleconferencing, gaming, movie animation, or as a virtual assistant.
@@ 3.2 NPHM
NPHM [18] is a morphable head model that generates a signed distance field (SDF) of a human head given an identity code $z_{id}$ and an expression code $z_{exp}$. We employ COLMAP [63, 62] to obtain pointclouds $\{\mathcal{P}_{t}\}$ for each timestep $t$ of a multi-view video sequence of the person. Subsequently, we fit NPHM to each pointcloud $\mathcal{P}_{t}$ to obtain $z_{id}$ and $z^{t}_{exp}$. We use a variant of NPHM, namely MonoNPHM [19], that uses a backward deformation field instead of the originally proposed forward deformation field as it simplifies the fitting process. Formally, we obtain the NPHM fitting as follows:
@@ 3.6 Deferred Diffusion
NeRFace [16] DNR [72] MVP [44] DiffusionRig [12] Ours GT
Figure 3: Qualitative results for self-reenactment. We compare against 3D methods (NeRFace [16], Mixture of Volumetric Primitives [44]) and methods employing 2D renderers (Deferred Neural Rendering [72], DiffusionRig [12]). Note that the slightly gray background for DiffusionRig is caused by their training scheme using $\epsilon$-prediction (see Sec. 3.1). Our method consistently produces more expressive facial performances while simultaneously providing more detailed renderings.
Source Actor NeRFace [16] DNR [72] DNR + GAN DiffusionRig [12] Ours
@@ 4.1 Training
For our experiments, we employ the recently released multi-view video dataset from NeRSemble [36]. The dataset contains 26 different multi-view sequences per person captured with 16 cameras, featuring different facial expressions, sentences and emotions. We use BackgroundMattingV2 [41] to remove the background and a segmentation network to remove the torso [91]. We fit NPHM on the video sequences of 8 different persons to create pairings of multi-view videos and fitted meshes. In total, we use about 3300 timesteps per person. We train one DiffusionAvatar for each person on these pairings with Adam [35] using a learning rate of $41e-$ 4 for ControlNet and the expression conditioning layers, and a learning rate of $21e-$ 2 for the TriPlane feature lookups. We use a batch size of $B=8$ and spatial resolutions $H=W=512$. We employ random cropping and resizing of the target images to avoid overfitting. The size of our spatial feature maps is $512\times 512\times 16$ for each TriPlane and the ambient map. In total, this yields $16\times(3+1)+5+1+3=73$ channels for the input renderings $R_{t}$ fed to our diffusion-based renderer. The pre-trained LDM is a Stable Diffusion v2.1 network [59]. Training takes roughly two days on a single RTX A6000 for 100k gradient descent steps.
[…]
NeRFace [16] 23.0 0.279 6.76 5.37 1.06 0.053 0.787
DiffusionRig [12] 19.6 0.220 6.41 2.74 0.55 0.029 0.887
DNR [72] 24.5 0.226 7.32 2.06 0.63 0.027 0.903
DNR+GAN [72] 23.0 0.114 7.08 2.14 0.69 0.028 0.868
MVP [44] 23.6 0.221 7.02 3.42 0.78 0.034 0.882
Ours 24.9 0.081 7.55 1.79 0.50 0.023 0.917
Table 1: Quantitative Comparison: We report metrics for a self-reenactment scenario on unseen expressions and unseen views averaged over 8 persons from the NeRSemble dataset [36].
@@ 4.2 Experiment Setup
To measure the quality of a 3D avatar, we conduct experiments on two tasks: Self-reenactment and avatar animation. For self-reenactment, we hold out 2 complete sequences and the frontal camera of every sequence during training to measure both image quality and view consistency. During testing, we provide the holdout fitted meshes of the same person and render from the novel view. For avatar animation, we first generate NPHM meshes using the same identity code $z_{id}$ as the avatar and a sequence of expression codes $\{z^{t}_{exp}\}$ from a different person. We then render a video of the avatar from a novel view. To evaluate the realism of this expression transfer, we conduct a user study where we show the animated avatars to participants.
@@
We employ two paired-image metrics to measure the quality of individual generated frames: Peak Signal-to-Noise Ratio (PSNR) and Learned Perceptual Image Patch Similarity (LPIPS) [90]. Furthermore, we measure visual similarity of a rendered video to its ground truth counterpart with the perceptual video metric JOD [47]. Finally, we make use of several face-specific metrics: Average Keypoint Distance (AKD) measured in pixels with keypoints estimated from PIPNet [30], and cosine similarity (CSIM) of identity embeddings based on ArcFace [10]. We further follow [9] and utilize DECA [14] to measure Average Expression Distance (AED) and Average Pose Distance (APD) between the rendered and the ground truth images.
@@
Mixture of Volumetric Primitives (MVP) [44]. An encoder-decoder architecture that generates a set of small feature grids rigged to a template mesh. We provide a FLAME tracking to learn the driving.
Deferred Neural Rendering (DNR) [72]. A screen-space decoder built on top of Pix2Pix [94] that uses learnable neural textures rigged to a tracked Basel Face Model [55]. We provide our fitted NPHM meshes as input for better comparison. Additionally, we compare against a version of DNR that also uses an adversarial loss for sharper images that we call DNR+GAN in the following.
@@ 4.3 Self-Reenactment
In Tab. 1, we compare DiffusionAvatars to other state-of-the-art approaches on the self-reenactment task. Results are averaged over avatars generated from eight different persons. Our method shows the overall best performance with the biggest improvement in perceptual and face-specific metrics. The results indicate that DiffusionAvatars generates sharper images (LPIPS) that are more temporally consistent (JOD). Furthermore, our method better maintains the person’s identity when rendered from a novel view (CSIM) than the baselines. Finally, facial features such as the position of mouth corners and eyebrows synthesized by our method match the ground truth expressions more closely (AKD, AED, APD).
@@
NeRFace DiffusionRig DNR DNR+GAN Ours
@@
[…]
Table 2: User Study: For our user study on avatar animation, we collect 735 responses from 49 participants to measure Visual Quality (VQ) and Driving Fidelity (DF) on a scale of 1-5.
@@ 4.4 Avatar Animation
Due to the task’s inherent lack of any ground truth, we perform a user study to validate avatar animation. We collect 735 responses from 49 participants on two questions:
(i) Visual Quality (VQ): What is the overall quality of the avatar animation?
(ii) Driving Fidelity (DF): How closely does the animated avatar match the expressions of the driving sequence?
We present two videos side-by-side: One sequence of a source actor performing various expressions, and a video of another person’s avatar with the expressions transferred to it. The results of our user study can be seen in Tab. 2. Our method achieves the highest scores for both visual quality and similarity to the driving sequence. The difference is particularly pronounced for the visual quality score. We attribute this to the fact that users heavily penalize any visual artifacts immediately obvious in video renderings. These perceptual differences in the task of avatar animation are also apparent in the qualitative comparison shown in Fig. 4.
@@ 4.5 Ablations
In Tab. 3, we examine the design choices of our method. All ablation experiments are conducted on 3 avatars on the self-reenactment task. We refer to our supplementary material for a qualitative comparison.
@@
Table 3: Ablation of architectural choices.
@@
To study whether diffusion is necessary, we train our architecture without the pre-trained LDM to directly predict the output in a single forward pass, similar to Deferred Neural Rendering. Such a method performs notably worse, especially in the LPIPS metric, indicating that diffusion is important for sharp renderings.
@@
We compare DiffusionAvatars with a version trained on renderings of a fitted FLAME instead of NPHM. Since NPHM models facial geometry as an SDF, it possesses a much higher representational power than FLAME. As a result, the renderings are much closer to the actual desired appearance of the head. This is also reflected in our ablation study, where using FLAME gives consistently worse performance. This is especially true for the AKD metric, indicating that our diffusion-based neural renderer cannot reenact certain complex expressions based on FLAME alone.
@@
Turning off expression conditioning worsens overall performance, presumably because the expression codes help to distinguish subtle expression details that can hardly be inferred from the NPHM mesh alone. It also helps DiffusionAvatars to synthesize expressions not captured by NPHM, such as tongue movement, by enabling the network to correlate such appearances with certain expression code combinations.
@@
We compare three spatial feature mapping approaches: (i) Our Triplanes, (ii) a simpler approach based on spherical feature mapping, and (iii) no spatial features at all. For spherical feature mapping, we create an approximate UV map for each NPHM mesh by projecting the 3D canonical coordinates for each vertex onto a sphere centered inside the head. Tab. 3 shows that using TriPlanes boosts most metrics, whereas the simpler spherical mapping brings almost no benefit.
@@ Acknowledgements
Appendix A Additional Results
We provide additional qualitative results for the self-reenactment task on three more people from the NeRSemble [36] dataset in Fig. 11.
A.1 Novel View Synthesis Results
@@
Table 4: Quantitative Comparison: We report metrics for novel-view-synthesis averaged over five persons from the NeRSemble dataset.
A.2 Ablation of Number of Training Views
We analyze the data efficiency of DiffusionAvatars on one person of the Multiface dataset [80] in Fig. 5. Our method only suffers negligibly from the removal of train views, as opposed to MVP. We attribute this to NPHM’s strong prior over 3D head shapes which allows fitting to sparse pointclouds, while direct 3D methods like MVP struggle more to recover accurate surfaces from sparse views.
Figure 5: Ablation of number of cameras on the Multiface dataset.
A.3 Qualitative Comparison of Ablations
FLAME [38] Ours GT
w/o exp. cond. Ours GT
w/o LDM prior Ours GT
A.4 Experiments on the Multiface dataset
We conduct additional experiments on 16 cameras from the Multiface dataset [80] in Tab. 5. Although MVP has the advantage of a high-end textured mesh tracking pre-computed from 100+ cameras there, DiffusionAvatars performs competitively showing that our method (i) can compete with recent 3D methods and (ii) generalizes to different datasets. In Fig. 9, we show that our method produces sharper renderings than MVP on the Multiface dataset.
[…]
MVP 26.4 0.126 7.92 2.25 0.655 0.026 0.904
Ours 25.5 0.077 7.77 1.86 0.503 0.025 0.914
Table 5: Quantitative comparison for self-reenactment averaged over 2 persons from the Multiface dataset.
MVP [44] Ours GT
Figure 9: Qualitative comparison for self-reenactment on the Multiface dataset.
@@ Appendix B User Study Setup
DNR [72] DNR+GAN MVP [44] DiffusionRig [12] Ours GT
Figure 11: Qualitative results for self-reenactment.
Appendix C Analysis of Temporal Consistency
Since DiffusionAvatars is based on 2D diffusion models, the generated images may be susceptible to view-inconsistencies and screen-space stitching artifacts. Following StyleGAN3 [32], we therefore perform an analysis of the temporal consistency of generated videos. In figure Fig. 12, we compare our method to MVP [44] and Deferred Neural Rendering [72]. Our finding is that DiffusionAvatars generates images with more high-frequency detail than 3D methods while being significantly less prone to texture-sticking than other 2D methods.
MVP DNR+GAN Ours GT
$\leftarrow$ time $\rightarrow$ $\leftarrow$ time $\rightarrow$ $\leftarrow$ time $\rightarrow$ $\leftarrow$ time $\rightarrow$
Figure 12: Analysis of Temporal Consistency. We extract a short vertical segment of pixels from each frame of a generated video and concatenate them left-to-right (bottom). The desired result is that texture details move together with the underlying head geometry. It can be clearly seen that combining Deferred Neural Rendering with an adversarial loss (DNR+GAN) causes severe texture sticking as the person rotates their head to the left. On the other hand, MVP and our method do not suffer from this issue. Additionally, as opposed to MVP, our method generates much of the high-frequency detail that is also present in the ground truth video.
Appendix D Model Architecture
Fig. 13 depicts DiffusionAvatars’s model architecture.
Appendix E Societal Impact
DiffusionAvatars provides the means to generate realistic images of faces with control over viewpoint, pose, and expression. As such, an avatar could be misused to the original person’s disadvantage. However, we focus on 3D applications such as immersive teleconferencing or character animation, not indistinguishable 2D video synthesis. Our method does not model lighting, the background, or the torso. Therefore, we believe a detection system can be built to identify generations of our method by exploiting such characteristics [60].
@@ Appendix E Societal Impact
(a) Diffusion U-Net (b) Expression Attention Block
Figure 13: Architecture of DiffusionAvatars. We build on the U-Net architecture of a pre-trained Stable Diffusion [59] model. While we keep all original layers fixed, we introduce an additional cross attention layer in each block that allows spatial features to attend to expression tokens extracted from NPHM’s expression codes. That way, subtle expression details can influence the image formation process. Additionally, following the ControlNet [89] paradigm, the encoder part of the diffusion U-Net is copied and conditioned on the NPHM rasterizations. The feature maps produced by this ControlNet are added to the existing skip connections of the U-Net, allowing the diffusion model to generate articulated facial images while keeping the knowledge from its pre-training.